OpenStudy (anonymous):

you have a piggy bank containing 78 coins in dimes and quarters. If the piggy bank contains $12.60, how many dimes in the piggy bank?

OpenStudy (anonymous):

so if q = quarters and d = dimes

OpenStudy (anonymous):

q+d=78

OpenStudy (anonymous):

.25q+.10d=$12.60

OpenStudy (baldymcgee6):

that's perfect, now just solve the system

OpenStudy (anonymous):

ok so i would multiply 4 through out equation 2 right?

OpenStudy (baldymcgee6):

personally, I would use substitution instead

OpenStudy (anonymous):

like .25q=$12.60-10d?

OpenStudy (baldymcgee6):

q+d=78 q= 78-d 0.25q+0.10d=$12.60 0.25(78-d) + 0.10d = 12.60

OpenStudy (baldymcgee6):

the first equation is much easier to solve for one variable, then sub into the other.

OpenStudy (anonymous):

19.5-0.25d+0.10d=12.60 turns into 19.5+.15d=12.60 and then .15d=6.9?

OpenStudy (baldymcgee6):

so how many dimes were in the bank? then we know from before that q= 78-d

OpenStudy (anonymous):

46 dime and 15 q?

OpenStudy (baldymcgee6):

your quarters is wrong, use q= 78-d

OpenStudy (anonymous):

oh thought it was suppose to be 61 coins my mistake but then q=32?

OpenStudy (baldymcgee6):

right, good job.

OpenStudy (anonymous):

wow thanks the way teacher taught me was confusing is substitution easier in these situations?

OpenStudy (baldymcgee6):

when you are dealing with decimals, and small number, substitution is generally easier as you cannot always find a convenient multiple